# Complexity of deep computations via topology of function spaces

### v3

@@ Abstract.
We use topological methods to study the complexity of deep computations and limit computations. We use topology of function spaces, specifically, the classification of Rosenthal compacta, to identify new complexity classes. We use the language of model theory, specifically, the concept of independence from Shelah’s classification theory, to translate between topology and computation. We use the theory of Rosenthal compacta to characterize approximability of deep computations, both deterministically and probabilistically.
@@ Introduction
Informed by model theory, to each computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model. This allows us to view computations in any given computational model as elements of a space of real-valued functions, which is called the space of types of the model. The idea of embedding models of theories into their type spaces is central in model theory. In the context of this paper, the embedding of computations into spaces of types allows us to utilize the vast theory of topology of function spaces, known as $C_{p}$-theory, to obtain results about complexity of topological limits of computations. As we shall indicate next, recent classification results for spaces of functions provide an elegant and powerful machinery to classify computations according to their levels of “tameness” or “wildness”, with the former corresponding roughly to polynomial approximability and the latter to exponential approximability. The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [TOD99]; in logic, the classification of theories developed by Shelah [SHE90]; and in statistical learning, the notions of PAC learning and VC dimension pioneered by Vapnik and Chervonenkis [VC74, VC71].
In a previous paper [ADI+24], the authors introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]). There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
@@
Deep computations arise as topological limits of standard (continuous) computations. Simplest among these is when the limit is sequential. In topology, the first Baire class, or Baire class 1 consists of functions (also called simply “Baire-1”) arising as pointwise limits of sequences of continuous functions. Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity above the Baire class 0 which (by definition) consists of continuous functions.
@@
In this paper, we apply the theory of spaces of Baire class 1 functions to identify and separate new complexity classes of deep computations. We obtain a new “tame vs wild” dichotomy for deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and further, a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
@@
Todorčević’s trichotomy concerns Rosenthal compacta, which are topological spaces that fit our context superbly: they are defined as compact spaces that are homeomorphic to subspaces of the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence. Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways. Since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces. Todorčević’s trichotomy for Rosenthal compacta has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22]. It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory. At the time of writing this paper, decades after his original argument, no elementary proof has been found [TOD23, HT19].
@@
Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of Non Independence Property (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concepts VC-dimension and Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory [VC71, VC74, VAL84].
@@
Refining Todorčević’s trichotomy, we invoke a more recent heptachotomy for separable Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; these authors identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
@@
In section 1, we present basic topological and combinatorial preliminaries, and in section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed). Section 3 is devoted to the classification of deep computations.
Throughout the paper, our results pertain to classical models of computation (particularly computations involving real-valued quantities that are known and manipulated to a finite degree of precision). The final section, Section 4, introduces a probabilistic viewpoint, the development of which we intend to pursue in future research, extending the present framework to encompass non-deterministic and quantum computations.
1. General topological preliminaries: From continuity to Baire class 1
In this section we present some preliminaries from general topology and function space theory. In the interest of completeness, we include some proofs that may be safely skipped by readers familiar with these topics.
Recall that a subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets. A space is metrizable if its topology agrees with the topology induced by some metric therein. Two such metrics inducing the same topology may induce quite different properties in the category of metric spaces. For example, the interval $(0,1)$ with the usual metric (as a subset) of the reals is not complete; however, $(0,1)$ is homeomorphic to the real line, which is complete with respect to the usual metric thereon. In a metrizable space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
A Polish space is a separable and completely metrizable topological space, i.e., admitting some complete metric inducing its topology. Although other (possibly incomplete) metrics may induce the same topology, being Polish is a purely topological property. One of the most important Polish spaces is the real line $\mathbb{R}$; the others include the Cantor space $2^{\mathbb{N}}$ and the Baire space $\mathbb{N}^{\mathbb{N}}$. The class of Polish spaces is closed under countable topological products; in particular, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology), and the space $\mathbb{R}^{\mathbb{N}}$ of sequences of real numbers are all Polish. Recall that the product topology on these spaces is the topology of pointwise convergence: a sequence converges in the space if and only if it converges at each coordinate index.
@@ 1. General topological preliminaries: From continuity to Baire class 1
A subset of a Polish space is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set. In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
For a proof, see [ENG89, 4.3.24].
Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence. The space $C_{p}(X,\mathbb{R})$ of continuous real functions on $X$ is denoted simply $C_{p}(X)$. A natural question is, how do topological properties of $X$ translate into $C_{p}(X)$ and vice versa? This general question, and the study of these spaces in general, is the concern of $C_{p}$-theory, an active field of research in general topology which was pioneered by A. V. Arhangel’skiĭ and his students in the 1970’s and 1980’s [ARK92]. This field has found many applications in model theory and functional analysis. For a recent survey, see [TKA11].
A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions. In symbols, $f:X\to Y$ is Baire class 1 if there is a sequence of continuous functions $f_{n}:X\to Y$ such that for all $x\in X$, $f(x)=\lim_{n\to\infty}f_{n}(x)$. If $X$ and $Y$ are topological spaces, the space of Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$ (as above, $B_{1}(X,\mathbb{R})$ is denoted $B_{1}(X)$). Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)\subseteq Y^{X}$ and we give these the topology (called the topology of pointwise convergence) inherited from the product topology of $Y^{X}$. The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles. His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits. An elementary fact about Baire class 1 functions is that they are continuous except on a set of first category (also called a meager set, a set of first cateogory is the countable union of sets whose closure has empty interior; intuitively, these sets are “topologically small”). Thus, Baire class 1 functions are continuous on a “topologically large” subset of their domain.
A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $F_{\sigma}$). Every metrizable space (hence, every Polish space) is perfectly normal.
A topological space $X$ is Baire if every countable intersection of dense open sets is dense. The Baire Category Theorem states that every compact Hausdorff or completely metrizable space (hence, every Polish space) is Baire.
@@
A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact. Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest in analysis and topological dynamics. We begin with the following well-known result. Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$. We include a proof for the reader’s convenience:
@@
Let $X$ be a Polish space and $A\subseteq B_{1}(X)$ be pointwise bounded. The following are equivalent:
@@
    AA is relatively compact in $B_{1}(X)$.
@@
    AA is relatively countably compact in $B_{1}(X)$, i.e., every countable subset of $A$ has a limit point in $B_{1}(X)$.
@@
    $\overline{A}\subseteq B_{1}(X)$, where $\overline{A}$ denotes the closure in $\mathbb{R}^{X}$.
@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
In metrizable spaces, points of the closure of some subset can always be approximated by points in the set proper, via a convergent sequence. For more complicated spaces, such as $C_{p}$-spaces, this fails in remarkable ways. The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ has no convergent subsequences, in $\mathbb{R}^{X}$. In a sense, this example exhibits the worst failure of sequential convergence possible. The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) is homeomorphic to the Stone-Čech compactification of the discrete space of natural numbers, usually denoted $\beta\mathbb{N}$, which is an important object of study in general topology.
@@
If $X$ is Polish and $\{f_{n}\}\subseteq C_{p}(X)$ is pointwise bounded, then $\{f_{n}\}_{n\in\mathbb{N}}$ has a convergent subsequence, or a subsequence whose closure in $\mathbb{R}^{X}$ is homeomorphic to $\beta\mathbb{N}$.
Rosenthal’s Dichotomy states that a pointwise bounded set of continuous functions contains either a convergent subsequence, or a subsequence whose closure is essentially the same as the example mentioned in the previous paragraphs (i.e., “wildest” possible). The genesis of this theorem was Rosenthal’s $\ell_{1}$-Theorem”, which states that a Banach space includes an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences), or else every bounded sequence therein is weakly Cauchy. The $\ell_{1}$-Theorem connects diverse areas: Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, a dichotomy paralleling the $\ell_{1}$-Theorem holds: Either every point of the closure of a set of functions is a Baire class 1 function, or there is a sequence in the set behaving in the wildest possible way. This result is usually not phrased as a dichotomy, but rather as an equivalence as in Theorem 1.5 below.
First, we introduce some useful notation. For any set $A\subseteq\mathbb{R}^{X}$ and any real $a$, define
@@
(In case $A=\emptyset$, we define $X^{\emptyset}_{\geq a}=X=X^{\emptyset}_{\leq a}$.) For any sequence $\{f_{n}\}\subseteq\mathbb{R}^{X}$ and $I\subseteq\mathbf{N}$, define $I^{\complement}\coloneqq\mathbb{N}\setminus I$ and $f_{I}\coloneqq\{f_{i}:i\in I\}$.
@@
    For every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{I}}\cap X_{\geq b}^{f_{I^{\complement}}}=\emptyset$.
@@
We say that a function collection $A\subseteq\mathbb{R}^{X}$ has the finitary No-Independence Property (NIP) if, for all sequences $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$, there exist finite disjoint sets $E,F\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}=\emptyset$. We say that such $E,F$ witness finitary NIP for $A$, $\{f_{n}\}$ and $a,b$.
A set $A\subseteq\mathbb{R}^{X}$ has the finitary Independence Property (IP) if it does not have finitary NIP, i.e., if there exists a sequence $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$ such that for every pair of finite disjoint sets $E,F\subseteq\mathbb{N}$, we have $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}\neq\emptyset$.
@@
If we insist on witnesses $E,F\subseteq\mathbb{N}$ such that $F=E^{\complement}$, we call the respective properties “BFT-NIP” (even weaker than countable NIP) and “BFT-IP” (even stronger than countable IP). Thus, Theorem 1.5 becomes that statement, for pointwise bounded function collections $A\subseteq C_{p}(X)$, that $A$ is relatively compact in $B_{1}(X)$ if and only if $A$ has BFT-NIP.
@@
Trivially (as per the preceding discussion), finitary NIP implies BFT-NIP. Reciprocally, assume that $X$ is compact and has finitary IP. Fix $A\subseteq C_{p}(X)$, a sequence $\{f_{n}\}\subseteq A$ and reals $r<s$. For any $I,J\subseteq\mathbb{N}$ (almost disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$. For $I\subseteq I^{\prime}\subseteq\mathbb{N}$ and $J\subseteq J^{\prime}\subseteq\mathbb{N}$, we have $X_{I,J}\supseteq X_{I^{\prime},J^{\prime}}$; moreover, $X_{I,J}=\bigcap_{E\subseteq I,F\subseteq J}X_{E,F}$, where the index variables $E\subseteq I$, $F\subseteq J$ range over finite subsets of $I,J$, respectively. Clearly, $E,F\subseteq\mathbb{N}$ witness finitary NIP for $\{f_{n}\}$ if and only if $X^{E,F}=\emptyset$.
Fix $I\subseteq\mathbb{N}$. Since $\{f_{n}\}\subseteq A\subseteq C_{p}(X)$ is a sequence of continuous functions, and $X$ is compact, the nested family $\{X_{E,F}:E\subseteq I,F\subseteq I^{\complement}\}$ consists of closed, thus compact, sets. Since $A$ has finitary IP by hypothesis, the nested family consists of nonempty sets, hence its intersection $X_{I,I^{\complement}}\neq\emptyset$ by compactness. This holds for arbitrary $\{f_{n}\}\subseteq A$ and $r<s$, so $A$ has BFT-IP. ∎
@@
Theorem 1.8 may be stated as the following dichotomy (under the assumptions): either $A$ is relatively compact in $B_{1}(X)$, or $A$ has IP (in either sense).

### v4

@@ Abstract.
We use topological methods and a suitable framework to study the complexity of limit computations (such as so-called “deep computations” implied by recurrent neural networks). By using topology of function spaces —specifically, the classification of Rosenthal compacta— we identify new computational complexity classes. The language of model theory, particularly, the concept of independence from Shelah’s classification theory, provides a medium to translate between topology and computation. The theory of Rosenthal compacta allows us to characterize approximability of deep computations, both deterministically and probabilistically.
@@ Introduction
The main result of the paper is the identification of a hierarchy of four distinct complexity classes and seven minimal prototypes for limit computations. This is obtained by invoking results from topology of spaces of functions, namely, Todorčević’s trichotomy for Rosenthal compacta [TOD99] and the Argyros-Dodos-Kanellopoulos heptachotomy [ADK08], and transferring them into the realm of computation. The medium that we use to translate between topology and computation is model theory, specifically, Shelah’s classification theory [SHE90].
The translation between topology and computation proceeds as follows. To every computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model. This association allows us to view computations in any given computational model as elements of a space of real-valued functions, which is called the space of types of the model. The idea of embedding models of theories into their type spaces is borrowed from model theory, where types spaces play a central role. The embedding of computations into spaces of types then allows us to utilize the vast theory of topology of function spaces, known as $C_{p}$-theory, to obtain results about complexity of topological limits of computations. As we shall indicate next, recent classification results for spaces of functions provide an elegant and powerful machinery to classify computations according to their levels of “tameness” or “wildness”, with the former corresponding roughly to polynomial approximability and the latter to exponential approximability. The model-theoretic viewpoint of spaces of types thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: in topology, the classification of Rosenthal compacta pioneered by Todorčević [TOD99]; in logic, the classification of theories developed by Shelah [SHE90]; and in statistical learning, the notions of PAC learning and VC dimension pioneered by Vapnik and Chervonenkis [VC74, VC71].
In a previous paper [ADI+24], the authors introduced the concept of limits of computations, which they called ultracomputations (given that they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]). There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use primarily the term “deep computation”.
@@
In this paper, we refine the dichotomy of [ADI+24], and obtain new complexity classes for deep computation. The “wild” part of the dichotomy from [ADI+24] is further subdivided between “tame” and “wild” classes; three new levels of tameness are identified within this new subdivision, and seven distinct minimal prototypes of tameness are detected.
The finer classification results of this paper are obtained through more contemporary topological and model-theoretic tools than in the previous paper [ADI+24]. From the topological side, instead of Grothendieck’s classical criterion, we use the modern theory of Rosenthal compacta, and from the model theory side, instead of stability, we use neostability. The main tool imported from neostability is Shelah’s Independence Property, or rather, its negation, which is commonly referred to as “NIP”. More on this below.
@@
Deep computations arise as topological limits of standard (continuous) computations. In topology, the first Baire class, or Baire class 1 consists of functions (also called simply Baire-1) arising as pointwise limits of sequences of continuous functions. Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity above the Baire class 0, which (by definition) consists of continuous functions.
@@
We apply a result of Bourgain, Fremlin and Talagrand from the late 1970s on compact subsets of the first Baire class to obtain a new “tame vs wild” dichotomy for deep computations [BFT78]; then, we apply a trichotomy of Todorčević, from the late 90s, for Rosenthal compacta to identify new tame classes [TOD99].
@@
A Rosenthal compactum is a compact topological space that is homeomorphic to a subspace of the space of Baire class 1 functions on some Polish (i.e., separable, completely metrizable) space equipped with the topology of pointwise convergence. Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways. Since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces. Todorčević’s trichotomy for Rosenthal compacta has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22]. It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory. At the time of writing this paper, decades after his original argument, there are no elementary proofs of the trichotomy; in fact, Todorčević’s methods have been analyzed and exploited further [TOD23, HT19].
@@
Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of No-Independence Property (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concepts VC-dimension and Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory [VC71, VC74, VAL84].
@@
Beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for separable Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]. These authors identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes.
From the point of view of approximation theory, it is noteworthy that in the context of Grothendieck [GRO52] and Bourgain-Fremlin-Talagrand [BFT78], compactness and sequential compactness coincide; thus, the results of these papers allow us to ensure that deep computations, which in principle are abstract ultrafilter limits (over a possibly uncountable set), can always be approximated by concrete sequences of standard computations.
@@
In section 1, we present basic topological and combinatorial preliminaries. In section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed). Section 3 is devoted to the classification of deep computations.
Throughout the paper, our results pertain to classical models of computation (particularly computations involving real-valued quantities that are known and manipulated to a finite degree of precision). The final section, Section 4, introduces a probabilistic viewpoint, the development of which we intend to extend in future research.
1. Topological preliminaries:
From continuity to Baire class 1
In this section we recall notions and results from general topology and function space theory. For simplicity, we assume that all spaces are Hausdorff unless explicitly stated otherwise.
A subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets. In a metric space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
A Polish space is a separable and completely metrizable topological space, i.e., one admitting a complete metric inducing its topology. Although other (possibly incomplete) metrics may induce the same topology, being Polish is a purely topological property. The class of Polish spaces includes the real line $\mathbb{R}$ as an important example, and is closed under countable topological products. [^1] Recall that the product topology is the topology of pointwise convergence. In a product space, a sequence (or a net, or a filter) converges if and only each coordinate thereof converges. In particular, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology), and the space $\mathbb{R}^{\mathbb{N}}$ of sequences of real numbers are Polish spaces, each of considerable interest.
@@ 1. Topological preliminaries:
[ENG89, 4.3.24] A subset of a Polish space is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set. In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
A subset $A$ of a topological space is said to be of the first category if $A$ is included in a countable union of closed nowhere dense sets; otherwise, $A$ is of the second category. A first-class subset $A\subseteq X$ is also called a mager set; its complement $X\setminus A$ is comeager. A topological space $X$ is Baire if every comeager subset thereof is dense. In Baire spaces, a meager subset is “everywhere small” in the sense that it includes no nonempty open set. The Baire Category Theorem states that every compact Hausdorff or completely metrizable space (hence, every Polish space) is Baire. [^2] More generally, any Čech-complete space is Baire [ENG89, Theorem 3.9.3]. (All compact Hausdorff and all complete metric spaces are Čech-complete.)
Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence, i.e., with the relative (subspace) product topology $C_{p}(X,Y)\subseteq Y^{X}$—where $Y$ carries its topology, but $X$ serves merely as a (discrete) index set stripped of any topology. The space $C_{p}(X,\mathbb{R})$ of continuous real-valued functions on $X$ is denoted simply $C_{p}(X)$. A natural question is, how do topological properties of $X$ translate into $C_{p}(X)$ and vice versa? This general question, and the study of these spaces in general, is the concern of $C_{p}$-theory.
A function $f\in Y^{X}$ between topological spaces $X$, $Y$ is said to be of Baire class 1, or a Baire-1 function, if there is a sequence $(f_{n})_{n\in\mathbb{N}}\subseteq C_{p}(X,Y)$ such that $f=\lim_{n\to\infty}f_{n}$ (i.e., $f(x)=\lim_{n\to\infty}f_{n}(x)$ for all $x\in X$). An elementary argument shows that Baire-1 functions are continuous everywhere except on a set of the first category. If $X$ and $Y$ are topological spaces, the space of Baire-1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$; thus, $C_{p}(X,Y)\subseteq B_{1}(X,Y)\subseteq Y^{X}$ are topological inclusions. We abbreviate $B_{1}(X,\mathbb{R})$ as $B_{1}(X)$. Relatively compact subsets of $C_{p}(X)$ and of $B_{1}(X)$ (particularly for $X$ Polish) are of interest in analysis and topological dynamics.
A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ set (equivalently, every open subset of $X$ is an $F_{\sigma}$ set). Every metrizable space (hence, every Polish space) is perfectly normal.
@@
The Lemma below is well known, but we include a proof for the reader’s convenience.
@@
Let $X$ be a Polish space and $A\subseteq B_{1}(X)$ be pointwise bounded. [^3] A set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there exists $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$. The following are equivalent:
@@
    AA is relatively compact in $B_{1}(X)$, i.e., the closure of $A$ in $B_{1}(X)$ is compact;
@@
    AA is relatively countably compact in $B_{1}(X)$, i.e., every (infinite) countable subset of $A$ has a limit point in $B_{1}(X)$;
@@
    $\overline{A}\subseteq B_{1}(X)$, where $\overline{A}$ denotes the closure of $A$ in $\mathbb{R}^{X}$.
@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
In metrizable spaces, the closure of a subset consists of limits of convergent sequences therein. In general spaces, including many $C_{p}$-spaces, such a characterization of closures fails in remarkable ways.
The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$; however, $(p_{n})_{n\in\mathbb{N}}$ has no convergent subsequences in $\mathbb{R}^{X}$ ([TOD97, Chapter 1.1]). In a sense, this example exhibits the worst failure of sequential convergence possible: The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) is homeomorphic to the Stone-Čech compactification $\beta\mathbb{N}$ of the (discrete) space of natural numbers. The space $\beta\mathbb{N}$ has cardinality $2^{2^{\omega}}$ (largest among all separable topological spaces) and, effectively, is the set of ultrafilters on $\mathbb{N}$, which possesses complex and fascinating topological and combinatorial properties. [^4] Our earlier proof [ADI+24] of the existence of deep equilibria in compositional computational structures relies on a classical result on the existence of idempotent elements of $\beta\mathbb{N}$.
@@
If $X$ is Polish and $(f_{n})_{n\in\mathbb{N}}\subseteq C_{p}(X)$ is pointwise bounded, then $(f_{n})$ has a convergent subsequence, or a subsequence whose closure in $\mathbb{R}^{X}$ is homeomorphic to $\beta\mathbb{N}$.
The genesis of the dichotomy was Rosenthal’s $\ell_{1}$-Theorem”, which states that a Banach space includes an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences), or else every bounded sequence therein is weakly Cauchy. The $\ell_{1}$-Theorem connects diverse areas: Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, a dichotomy paralleling the $\ell_{1}$-Theorem holds: Either every point in the closure of a set of functions is a Baire class 1 function, or there is a sequence in the set behaving in the wildest possible way. This result is usually not phrased as a dichotomy in the above manner, but rather as an equivalence (see Theorem 1.5 below).
First, we introduce some useful notation. For any $f\in\mathbb{R}^{X}$ and any real $a$, define
$X^{f}_{\leq a}$ $\coloneqq f^{-1}(-\infty,a],$ $X^{f}_{\geq a}$ $\coloneqq f^{-1}[a,+\infty).$
For any set $A\subseteq\mathbb{R}^{X}$ and any real $a$, define (by a slight abuse of notation)
@@
(In case $A=\emptyset$, we define $X^{\emptyset}_{\geq a}=X=X^{\emptyset}_{\leq a}$.) For any sequence $\{f_{n}\}\subseteq\mathbb{R}^{X}$ and $I\subseteq\mathbb{N}$, define $I^{\complement}\coloneqq\mathbb{N}\setminus I$ and $f_{I}\coloneqq\{f_{i}:i\in I\}$.
@@
    For every $(f_{n})_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that
    $X_{\leq a}^{f_{I}}\cap X_{\geq b}^{f_{I^{\complement}}}=\emptyset.$
@@
We say that a collection of functions $A\subseteq\mathbb{R}^{X}$ has the finitary No-Independence Property (NIP) if, for all sequences $(f_{n})_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$, there exist finite disjoint sets $E,F\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}=\emptyset$. We say that such $E,F$ witness finitary NIP for $(f_{n})$ and $a,b$.
A set $A\subseteq\mathbb{R}^{X}$ has the finitary Independence Property (IP) if it does not have finitary NIP, i.e., if there exists a sequence $(f_{n})_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$ such that for every pair of finite disjoint sets $E,F\subseteq\mathbb{N}$, we have $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}\neq\emptyset$.
@@
If we insist on witnesses $E,F\subseteq\mathbb{N}$ such that $F=E^{\complement}$, we call the respective properties “BFT-NIP” and “BFT-IP”. Thus, Theorem 1.5 becomes (for pointwise bounded function collections $A\subseteq C_{p}(X)$) that $A$ is relatively compact in $B_{1}(X)$ if and only if $A$ has BFT-NIP.
@@
Trivially (as per the preceding discussion), finitary NIP implies BFT-NIP. Reciprocally, assume that $X$ is compact and $A\subseteq C_{p}(X)$ has finitary IP. Fix a sequence $(f_{n})\subseteq A$ and reals $a<b$ such that $X^{f_{E}}_{\leq a}\cap X^{f_{F}}_{\geq b}\neq\emptyset$ for every finite disjoint $E,F\subseteq\mathbb{N}$. Fix $I\subseteq\mathbb{N}$. Since $(f_{n})\subseteq A\subseteq C_{p}(X)$ is a sequence of continuous functions, the collection $\mathcal{F}_{I}=\{X^{f_{n}}_{\leq a}:n\in I\}\cup\{X^{f_{m}}_{\geq a}:m\in I^{\complement}\}$ consists of closed (non-empty) sets. Since all finite $E\subseteq I$ and $F\subseteq I^{\complement}$ witness finitary IP for $(f_{n})$ and $a,b$, we see that the collection $\mathcal{F}_{I}$ has the finite intersection property. By compactness, $X^{f_{I}}_{\leq a}\cap X^{f_{I^{\complement}}}_{\geq b}=\bigcap\mathcal{F}_{I}\neq\emptyset$. ∎
@@
Theorem 1.8 may be stated as the following dichotomy (under the assumptions) for any compact metrizable $X$: either $A$ is relatively compact in $B_{1}(X)$, or $A$ has IP (in either sense).